in a club with 12 members, how many ways can a slate of three officers consisting of president, vice president, and secretary/treasurer be chosen?
Answer #1

In this problem, there is 12 member in a club. One has to choose a total of 3 officers consisting President, Vice-president & Secretary.

After choosing a person, the total member in the club will decrease by 1 because we have selected 1 person

So for president, no. of ways= 12C1

So for Vice-President, no. of ways= 11C1

So for Secretary, no. of ways= 10C1

Hence, total no. of ways= 12C1 * 11C1 * 10C1= 12* 11* 10 = 1320